What is the probability that the mean weekly income of these 50 workers is less than 448 pesos?

1
Expert's answer
2022-05-30T03:36:30-0400

To calculate that, we should know the mean and the standard deviation. To find P we use formula:

"P(\=x<448)=P(z<\frac{448-\mu}{\sigma/\sqrt{n}})"

For example, we have "\mu=500, \sigma=160"

Then "P(z<\frac{448-\mu}{\sigma/\sqrt{n}})=P(z<\frac{448-500}{160/\sqrt{50}})=P(z<-2.29)=0.011"

(We found p using z-table).

So to calculate that you should substitute "\mu,\sigma" for your numbers and find P.